OpenStudy (anonymous):

Graph the function f(x) = (x + 3)3 by hand and describe the end behavior.

OpenStudy (anonymous):



OpenStudy (anonymous):

looks identical to \(y=x^3\) only shifted 3 units to the left if you know the graph of \(y=x^3\) then you can draw this one as well

OpenStudy (anonymous):

it is not a line

OpenStudy (anonymous):

y = (x+3)3 = 3x+9 Its a line, isn't it? @satellite73

OpenStudy (anonymous):

it is a line for sure

OpenStudy (anonymous):

A straight line with slope = 3 and y-intercept=9

OpenStudy (anonymous):

true, but i am assuming it is \(f(x)=(x+3)^3\) and not \((x+3)3\)

OpenStudy (anonymous):

Soo is it a line or not?

OpenStudy (anonymous):

it depends on whether it is \[f(x)=(x+3)\times 3=3x+9\] i.e. x plus 3 times 3 or \[f(x)=(x+3)^3\] i.e. x plus 3 cubed

OpenStudy (anonymous):

of course it is aline